# Explaining Context Length Scaling and Bounds for Language Models

### v2

@@ C.3 Synthetic dataset: Entropy in Intrinsic Space, and Entropy for output layer
For our synthetic dataset, if we view the Context Feature Vector shown in Figure 8 as the feature in the Intrinsic Space, then the best strategy for the context encoder is to generate the answer for all subtasks (it can see) in the Intrinsic Space (since it cannot see the task bits). This would lead to an Entropy of $H_{IS}=T\log 2$ in the Intrinsic Space.
@@
The entropy of the output layer is, however, $H_{output}=\log 2$ since the answer bits $0,1$ have the same probability. In this way, the answer of the output layer actually corresponds to one dimension in the Intrinsic Space, which should be the exact dimension at which the answer of the current task is stored. Therefore, $H_{output}=1/T*H_{IS}$, which explains why the Entropy for output logits is linear to the Entropy for Intrinsic Space.
@@
C.4 New Synthetic dataset: further experiments for Entropy
In the main paper, we use $T<ctl$, where the task number is smaller than the context bits. Here, we further study the case where $T>ctl$. Moreover, instead of letting tasks show at the same frequency, we use weighted frequency proportional to $max(bit_{1},bit_{2})^{-1.2}$ as the weights of tasks.
Here we train a single MLP (instead of multiple MLPs) as Bayes Model for different context lengths. The architecture is like:
Figure 12: New MLP trained with mask bits. For the example shown, the context length is $7$ and we mask the other task bits to be $0.5$.
@@ C.4 New Synthetic dataset: further experiments for Entropy
During training, $50\%$ of the samples are unmasked, while for the other $50\%$ samples, we mask the last $X$ task bits to be $0.5$, where $X$ is a random int from $60-10$ to $60-60$. This ensures our model to be able to handle mask bits, and also ensures it can learn uncommon tasks (relying on context bits that are at the end of the context bits) well.
The context bits MLP has hidden size $400$, and the prediction MLP has hidden size $300$. The linear embedding and the context feature have the same dimension $200$, and are added before being sent to the predict MLP. LeakyRELU with leaky hyperparameter $31e-$ 3 is used as an activation function for both MLPs. We train the model on a training set of $10000000$ and a validation set of size $1000000$, for $125$ epochs (and an early stopping setting of $25$ epochs, though the training process did not trigger early stopping).
To make sure that the trained model can be used to approximate the Bayes Model, we compare the model’s loss on validation set with context $ctl$ with the calculated minimum possible CE Loss for the task:
$MinCELoss(ctl)=\frac{\sum_{\text{task s.t. $max(bit_{1},bit_{2})>ctl$}}P(task)*\log 2}{\sum_{task}P(task)}$
And we obtain such a result:
@@
15 0.531 0.516
17 0.478 0.464
20 0.408 0.399
23 0.353 0.344
25 0.320 0.312
28 0.276 0.269
30 0.250 0.243
35 0.193 0.186
40 0.154 0.139
50 0.082 0.061
60 0.051 0.0
Table 1: Comparison between trained model and Bayes Model for Synthetic Data
We can see from Table 1 that the model is not too different from the Bayes Model: the BCE Loss only differs by around $0.02$. Thus, we can use the middle-layer-representation (shown as context feature in Figure 12) as the feature in Intrinsic Space to approximate the Bayes Model for $17\leq ctl\leq 50$.
Figure 13: Eigen value and CE results measured on trained model for Synthetic Dataset in this section. Left: relative eigen value vs. Index. Middle: Model CE Loss vs. $\sum_{i}\log rel\_eigval_{i}$ and Theoretical Min CE Loss vs. $\sum_{i}\log rel\_eigval_{i}$. Right: Correlation between minus Experimentally measured model CE Loss, minus Theoretically Minimum CE Loss, and $\sum_{i\leq N}\log rel\_eigval_{i}$ for different $N$s. All correlation factors are larger than $0.97$.

### v3

@@ B.5 Synthetic dataset: Entropy in Intrinsic Space, and Entropy for output layer
For our synthetic dataset, if we view the Context Feature Vector shown in Figure 5 as the feature in the Intrinsic Space, then the best strategy for the context encoder is to generate the answer for all subtasks (it can see) in the Intrinsic Space (since it cannot see the task bits). This would lead to an Entropy of $S=T\log 2$ in the Intrinsic Space.
@@
The entropy of the output layer is, however, $S_{output}=\log 2$ since the answer bits $0,1$ have the same probability. In this way, the answer of the output layer actually corresponds to one dimension in the Intrinsic Space, which should be the exact dimension at which the answer of the current task is stored. Therefore, $S_{output}=1/T*S_{IS}$, which explains why the Entropy for output logits is linear to the Entropy for Intrinsic Space.
@@
B.6 Generalized Synthetic dataset and Transformer-based Synthetic model: further experiments for Entropy
In the main paper, we set tasks to have same frequency. Here, we further study the case where tasks have different frequency. Moreover, we use another transformer and RoPE-based model to be closer to real-world cases. We use a $3$-layer causal Transformer, with embedding dimension $208$ and FFN dimension $832$, RoPE embedding with base frequency $4000$; input sequence length is always $60+1$, with $60$ context tokens (either 0, 1 or ?) and $1$ task tokens (chosen from task tokens of vocab size $200$).
Figure 10: New transformer and Rope-based model. Here, we use one task token to encode the task information.
@@ B.6 Generalized Synthetic dataset and Transformer-based Synthetic model: further experiments for Entropy
During training, $50\%$ of the samples are unmasked, while for the other $50\%$ samples, we mask the last $X$ task bits to be $0.5$, where $X$ is a random int from $60-10$ to $60-60$. This ensures our model to be able to handle mask bits, and also ensures it can learn uncommon tasks (relying on context bits that are at the end of the context bits) well. We train the model on a training set of $10000000$ and a validation set of size $1000000$, for $125$ epochs (and an early stopping setting of $25$ epochs, though the training process did not trigger early stopping).
To make sure that the trained model can be used to approximate the Bayes Model, we compare the model’s loss on validation set with context $ctl$ with the calculated minimum possible CE Loss for the task. As shown in Table 1 that the model is not too different from the Bayes Model: the BCE Loss only differs by around $0.001$. Thus, we can use the middle-representation (shown as context feature in Figure 10) as the feature in Intrinsic Space to approximate the Bayes Model for $17\leq ctl\leq 50$.
$MinCELoss(ctl)=(\sum_{\text{task s.t. $max(bit_{1},bit_{2})>ctl$}}freq(task)*\log 2)/{\sum_{task}freq(task)}$
@@
17 0.4648 0.4643
20 0.3988 0.3988
23 0.3438 0.3437
25 0.3119 0.3116
28 0.2687 0.2686
30 0.2429 0.2429
35 0.1867 0.1864
40 0.1390 0.1387
50 0.0613 0.0612
Table 1: Comparison between trained model and Bayes Model (minimum CE Loss) for Synthetic Data
Figure 11: Eigen value and CE results measured on trained model for Synthetic Dataset in this section. Upper figures: Eigen Value v.s. Index of Eigen Value. Lower-Left figure: Correlation between Cross Entropy Loss and Measured Intrinsic Dimension; Lower-Right figure: Correlation between Cross Entropy Loss and Measured Entropy. We see good linear relationship between CE Losses and Measured ID/Entropy from Lower figures.
@@
Appendix C Definitions of Intrinsic Space and Derived Properties
As mentioned in Section 2.1.2, in previous work[10, 12], as a common practice, the ‘Data Manifold’ is often defined as the middle feature representation of well-trained neural networks, and assumptions are made on this kind of mid-representation, with experiments to validate these assumptions. (Intrinsic Space is defined as the space where the Data Manifold lies.)
Meanwhile, the Data Manifold can be more formally defined by a mapping from input data to some Intrinsic Space which satisfies a certain set of properties, and mid-representation of well-trained neural networks are assumed to have such properties, which can be experimentally validated. These two perspectives are actually, equivalent to each other:
  • Perspective 1: Experiments show mid-representations of neural networks have certain properties $\rightarrow$ Data Manifold in Intrinsic Space satisfies such properties.
  • Perspective 2: Experiments show mid-representations of neural networks have certain properties $\rightarrow$ such mid-representation can be viewed as Data Manifold of Intrinsic Space that is defined to have such properties.
These two perspectives are equivalent to each other, and Perspective 1 is used in some previous work[10, 16].
In this section, we formally define the Intrinsic Space and formally derive related results, following Perspective 2.
C.1 Formal Definitions of Intrinsic Space
We define an intrinsic space to formalize the latent structure underlying natural language sequences. This space is independent of surface forms and aims to capture the semantic and syntactic essence of language contexts across different sequence lengths.
Setup.
Let $\mathcal{V}$ be a finite vocabulary and $\mathcal{X}=\mathcal{V}^{*}$ the set of all finite sequences over $\mathcal{V}$. Let $\mathcal{M}\subset\mathcal{X}$ denote the original data manifold of natural language, i.e., the support of the data distribution $p(x)$.
Definition.
An intrinsic space $\mathcal{Z}$ is a latent representation space defined by a mapping
$\Phi:\mathcal{X}_{\leq t}\to\mathcal{Z},$ (12)
where $\mathcal{X}_{\leq t}=\bigcup_{k=0}^{t}\mathcal{V}^{k}$ is the set of all language contexts of length $t$. The image of the original data manifold under this map is denoted $\mathcal{M}_{\mathcal{Z}}=\Phi(\mathcal{M}_{\leq t})\subset\mathcal{Z}$, or the data manifold (in Intrinsic Space). We require the following properties:
  • Predictive Consistency: There exists a decoder $\pi:\mathcal{Z}\to\Delta(\mathcal{V})$ such that
    $\pi(\Phi(x_{<t}))=p(x_{t}\mid x_{<t}),$ (13)
    i.e., the intrinsic representation enables accurate next-token prediction.
Moreover, there are some other properties assumed (separately) in our work, for which we give a formal definition here.
  • Uniform Information Gain: This assumption assumes the following linear relationship between predictive divergence and intrinsic dimension $dim(l)$:
    $D_{KL}(P,P_{l})=s\cdot\left(\dim(\infty)-\dim(l)\right)$ (14)
    for some constant $s>0$, which we interpret as the average number of bits of predictive information contributed by each intrinsic dimension. This is empirically observed in experiments.
  • Linear Entropy Relationship This assumption assumes that there exists constant $0<s<1,b$ and a sequence of tolerances $\{\varepsilon_{t}\}_{t\geq 0}$ with $\varepsilon_{t}\to 0$ such that for every context length $t$,
    $\Bigl{|}\,-s\;H\bigl{[}q_{t}(Z)\bigr{]}\;+\;b\;-\;H\bigl{[}p(\,\cdot\mid x_{<t})\bigr{]}\;\Bigr{|}\;\leq\;\varepsilon_{t}.$ (15)
    Where $q_{t}(\cdot)$ denotes probability density function. Equivalently, in the idealized zero-tolerance limit:
    $-s\;H\bigl{[}q_{t}(Z)\bigr{]}\;+\;b=\;H\bigl{[}p(x_{t}\mid x_{<t})\bigr{]}\quad\forall\,t.$ (16)
    It is worth mentioning that, we can easily derive the linear entropy relationship from the uniform information gain assumption, but not vice versa. Hence, linear entropy relationship is a weaker assumption compared to uniform information gain.
  • Lipschitz Differentiable Density This assumption assumes the density of data distribution is smooth in the intrinsic space:
    $\|\nabla q(z)\|\leq L$ (17)
    for some constant $L>0$
  • Finite $\epsilon$-negative moment: This assumption means the integral of the $\epsilon$-negative moment of the data distribution is finite:
[…]
Key Properties Assumed Derived Results
[10]Theorem 2 Bounded, Uniform Distribution, Data Scaling for Approx. Loss
Lipschitz Differentiable
Theorem 1, 2 in Appendix C.2 (Bounded,) Finite Negative Moment, Data Scaling for Approx. Loss
(corr. to Section 2.3) Lipschitz Differentiable
Theorem 3 in Appendix C.3 Predictive Consistency, Bayes Risk for Ntp of varied Ctl
(corr. to Section 2.2.1) Uniform Information Gain (Intrinsic Dimension perspective)
Theorem 4 in Appendix C.4 Predictive Consistency, Bayes Risk for Ntp of varied Ctl
(corr. to Appendix B) Linear Entropy Relationship (Information Entropy perspective)
Table 2: In this Section (Appendix C) we formulate results from previous sections with Theorems derived with defined assumptions and properties of intrinsic space in this section. Ntp refers to Next-token-prediction, Ctl refers to Context Length. We derive data scaling for approximation loss with weaker assumptions compared to [10], please refer to Theorem 1, 2 in Appendix C.2 for more details.
To conclude: if some space satisfies these properties, then the data representation is referred to as ‘Data Manifold’ in this space, and such properties would lead to further derivations in these work (including this work). In experiments, Middle-representation of neural networks are assumed (and shown) to have these kind of properties, hence explain some of the scaling behaviors they have.
C.2 Derivation for Data Scaling for Approximation Loss
Theorem 1 (Expected capped nearest–neighbour distance).
Let $\mathcal{Z}\subseteq\mathbb{R}^{d}$ ($d\geq 1$) and there exists a non-empty open set $U\in\mathbb{R}^{d}$ such that $U\subseteq\mathcal{Z}$ (i.e., $\mathcal{Z}$ is a $d$-dimensional region). Let $q:\mathcal{Z}\to[0,\infty)$ be a probability density satisfying
  1. 1.
    Lipschitz Differentiable:$\|\nabla q(z)\|\leq L$ for all $z\in\mathcal{Z}$;
  2. 2.
[…]
Draw i.i.d. samples $\mathcal{Z}_{D}=\{Z_{1},\dots,Z_{D}\}\sim q^{\otimes D}$ and define the capped nearest–neighbour distance
$R_{M}(Z_{i})\;=\;\min\!\bigl{\{}\,M,\;\min_{j\neq i}\|Z_{i}-Z_{j}\|\,\bigr{\}},\qquad M>0.$ (19)
Then, there exists constant $C=C(d,L,\epsilon,M,C_{\epsilon})$ and $D_{0}=D_{0}(d,L,\epsilon,M,C_{\epsilon})$ such that $\forall D>D_{0}$:
$\mathbb{E}_{\mathcal{Z}_{D}}\!\bigl{[}R_{M}(Z_{1})\bigr{]}\;\leq\;\begin{cases}C\,(\log D)^{\epsilon/(d+1)}\,D^{-\epsilon/(d+1)},&\text{if }\displaystyle\epsilon<\frac{d+1}{d},\\[6.0pt] C\,D^{-1/d},&\text{if }\displaystyle\epsilon\geq\frac{d+1}{d}.\end{cases}$ (20)
Thus, there exists constant $c=c(L,\epsilon,M,C_{\epsilon})$, such that:
$\mathbb{E}_{\mathcal{Z}_{D}}\!\bigl{[}R_{M}(Z_{1})\bigr{]}\;\leq\;C\;D^{-c/d}.$ (21)
Proof.
We write $Z_{1}$ for the distinguished point and $R(Z_{1})=\min_{j\neq 1}\|Z_{1}-Z_{j}\|$ for its exact nearest–neighbour distance, always capping by $M$ at the very end. Throughout the proof the expectation $\mathbb{E}[\cdot]$ is taken over the whole sample $\mathcal{Z}_{D}=(Z_{1},\dots,Z_{D})\sim q^{\otimes D}$.
Step 0. Notation.
$v_{d}:=\text{vol}(B_{1}(0)),\qquad c_{d}:=\frac{v_{d}}{2},\qquad r_{0}(z):=\frac{q(z)}{2L}.$
$v_{d}$ is the volume of a unit ball. Moreover, since $\|\nabla q\|\leq L$, whenever $r\leq r_{0}(z)$ one has $q(u)\geq\tfrac{1}{2}q(z)$ for every $u\in B_{r}(z)$.
Step 1. Exponential hole probability inside the Lipschitz ball.
Fix $z$ and $r\leq r_{0}(z)$. The mass of $q$ inside $B_{r}(z)$ satisfies
[…]
Conditioned on $Z_{1}=z$, the $(D-1)$ other points are i.i.d. $q$, so the conditional probability that all other points are sampled outside the ball $B_{r}(z)$ is:
$\Pr\!\bigl{(}R(z)>r\,\big{|}\,Z_{1}=z\bigr{)}=\bigl{(}1-\mu_{r}(z)\bigr{)}^{D-1}\;\leq\;\exp\!\bigl{[}-c_{d}(D-1)q(z)\,r^{d}\bigr{]}.$ (22)
Step 2. Density threshold and spatial split.
Define a data–dependent threshold
$\lambda_{D}\;:=\;\Bigl{(}\frac{2^{d+1}L^{d}d\log D}{c_{d}D}\Bigr{)}^{\!\frac{1}{d+1}}\quad(D\geq 2),\qquad\mathcal{H}_{D}:=\bigl{\{}z:q(z)\geq\lambda_{D}\bigr{\}},\quad\mathcal{L}_{D}:=\mathbb{R}^{d}\!\setminus\mathcal{H}_{D}.$
(The power $1/(d+1)$ is tuned to balance two error terms below.)
Step 3.1. Distribution in $\mathcal{H}_{D}$ (moderate or high density region).
For every $z\in\mathcal{H}_{D}$ set
$\rho(z,D)\;:=\;\Bigl{(}\frac{2d\log D}{c_{d}D\,q(z)}\Bigr{)}^{\!1/d}\;.$
Bound on $\rho$. Since $q(z)\geq\lambda_{D}$, $\rho(z,D)\leq r_{0}(z)$ and therefore (22) is valid for all $0<r\leq\rho(z,D)$.
Tail probability at $\rho$. With $r=\rho(z,D)$,
$\Pr\!\bigl{(}R(z)>\rho(z,D)\,\big{|}\,Z_{1}=z\bigr{)}\;\leq\;\exp[-2d\log D]=D^{-2d}.$
Because $R_{M}(Z_{1})=\min(R(Z_{1}),M)\leq M$,
$\mathbb{E}\bigl{[}R_{M}(Z_{1})\mathbf{1}_{\{R>\rho\}}\,\big{|}\,Z_{1}=z\bigr{]}\;\leq\;MD^{-2d}.$ (23)
Integral of $R$ up to $\rho$. Using (22),
[…]
Make the change of variable $t:=c_{d}(D-1)q(z)r^{d}$; then $r=(t/c_{d}(D-1)q(z))^{1/d}$ and $dr=\tfrac{1}{d}\,r\,dt/t$. The upper limit $r=\rho$ maps to $t=2d\log D$. Hence
[…]
Absorbing constants:
$\mathbb{E}\bigl{[}R(Z_{1})\wedge\rho(z,D)\,\big{|}\,Z_{1}=z\bigr{]}\;\leq\;C_{d,L}\,\bigl{(}Dq(z)\bigr{)}^{-1/d}.$ (24)
Average over $z\in\mathcal{H}_{D}$. Taking expectation over $Z_{1}$ first restricted to $\mathcal{H}_{D}$ and then combining (24) with (23),
$\mathbb{E}\bigl{[}R_{M}(Z_{1})\mathbf{1}_{\mathcal{H}_{D}}(Z_{1})\bigr{]}\leq C_{1}D^{-1/d}\;+\;MD^{-2d},\quad C_{1}:=C_{d,L}\bigl{(}\mathbb{E}[q(Z)^{-1/d}]\bigr{)}^{1/d}<\infty.$ (25)
Step 3.2. Distribution in $\mathcal{L}_{D}$ (ultra–low density region).
On $\mathcal{L}_{D}$ one has $q(z)^{\epsilon}\leq\lambda_{D}^{\epsilon}$, so by Hölder’s inequality
[…]
Since $R_{M}\leq M$,
$\mathbb{E}\!\bigl{[}R_{M}(Z_{1})\mathbf{1}_{\mathcal{L}_{D}}(Z_{1})\bigr{]}\leq MC_{\epsilon}\lambda_{D}^{\epsilon}=MC_{\epsilon}\bigl{(}\log D\bigr{)}^{\epsilon/(d+1)}D^{-\epsilon/(d+1)}.$ (26)
Step 4. Global bound.
Adding (25) and (26). For large enough $D$, the term $MD^{-2d}$ is higher-order small quantity compared to $D^{-1/d}$ or $D^{-\epsilon/(d+1)}$. Therefore,
$\mathbb{E}\bigl{[}R_{M}(Z_{1})\bigr{]}\;\leq\;C_{1}D^{-1/d}\;+\;MC_{\epsilon}(\log D)^{\epsilon/(d+1)}D^{-\epsilon/(d+1)}+o(\min(D^{-1/d},D^{-\epsilon/(d+1)})).$
Finally, compare the two powers of $D$. If $\epsilon<\frac{d+1}{d}$ then $\epsilon/(d+1)<1/d$ and the second term dominates; otherwise the first dominates. This yields the two–case estimate claimed. ∎
How large can $\epsilon$ be, for unbounded and bounded $\mathcal{Z}$?
  • It is usually assumed [10, 16] that $\mathcal{Z}$ is bounded: this assumption makes sense since in usual cases we approximate Intrinsic Space with middle feature representation of neural networks, which can indeed be bounded. For bounded $\mathcal{Z}$, it is possible for $\epsilon$ to be larger than $1+1/d$. We would like to mention that in [10], a constant distribution $q(z)=Const$ is assumed, where $\epsilon$ can be arbitrarily large and dominant rate $D^{-1/d}$ is derived: this is a much stronger assumption compared to Finite $\epsilon$-negative moment we assumed in our work.
[…]
  • For unbounded $\mathcal{Z}$, $\epsilon<1$ is the usual case; at $\epsilon=1$ the condition becomes $\int q^{0}=\text{Leb}(\operatorname{supp}q)<\infty$, i.e. compact support of finite measure. For most unbounded densities (Gaussians, sub-exponential, power-law) one only has $\epsilon<1$.
  • The comparison threshold $\frac{d+1}{d}$ is always $>1$ when $d\geq 1$; hence the dominant rate is
    $\begin{cases}D^{-\epsilon/(d+1)}&\text{for every admissible }1/d<\epsilon<1,\\ D^{-1/d}&\text{only if the support is compact and $\epsilon>1+1/d$.}\end{cases}$ (27)
    Thus $\epsilon$ can never “reach” the critical value $\tfrac{d+1}{d}$ unless $q$ is essentially bounded below on its support.
From Nearest-Neighbour Distance to Approximation Loss
  • Capped nearest-neighbour distance can be derived natually if one assume the maximum distance of neighboring points to be bounded by constant, or if one assume the Intrinsic Space $\mathcal{Z}$ is bounded.
  • Restate of Theorem 2 in [10]: Assuming $l(f),f,F$ be Lipschitz with contants $K_{L},K_{f},K_{F}$ and $l(F)=0$, $D$ be training dataset of size $D$ sampled i.i.d from $M_{d}$. Let $f(x)=F(x)\forall x\in D$. Then, for each training point $x$, let $\hat{x}$ be the nearest neighboring training data point, we have $L(D)\leq K_{L}(K_{f}+K_{F})\mathbb{E}_{D,x}[|x-\hat{x}|]$.
  • Combining Theorem 2 in [10] and previous results (nearest neighbour distance in this Appendix C.2), since Approximation loss of context length $l$ is $D_{KL}(P_{l},Q)$ which can be 00 when $Q=P_{l}$, thus satisfying the assumption of Theorem 2 in [10]. Thus, $L_{Approx}=C_{0}+A(l)/D^{c/dim}=C_{0}+A(l)/D^{\alpha(l)}$.
Therefore, we have:
Theorem 2 (Data Scaling for Approximation Loss).
Let $\mathcal{Z}\subseteq\mathbb{R}^{d}$ ($d>1$) be a d-dimensional region (exists non-empty open set $U\in\mathbb{R}^{d}$ such that $U\subseteq\mathcal{Z}$). $q:\mathcal{Z}\rightarrow[0,\infty)$ be probability density function satisfying Lipschitz Differentiable and Finite $\epsilon$-negative moment. Let $g:\mathcal{Z}\rightarrow P_{V}$ be a decoding mapping from intrinsic space $\mathcal{Z}$ to a distribution of tokens in vocabulary $V$, and $l(P_{V1},P_{V2})$ be KL divergence loss function (thus $l$ is zero for identical distributions). Assume $l(g(z_{1}),g(z_{2}))$ is differentiable and Lipschitz smooth for $z_{1}$ and $z_{2}$ with Lipschitz coefficient $L_{l}$.
Then, draw i.i.d. samples $\mathcal{Z}_{D}=\{Z_{1},\ldots,Z_{D}\}\sim q^{\otimes D}$, if $\min_{j\neq i}||\mathcal{Z}_{i}-Z_{j}||$ is bounded by $M$, then there exists constant $C=C(d,L,\epsilon,M,C_{\epsilon},L_{l})$ and $c=c(L,\epsilon,M,C_{\epsilon})$ such that for $D>D_{0}(d,L,\epsilon,M,C_{\epsilon})$:
$\min_{j\neq i}l(g(Z_{i}),g(Z_{j}))\leq C\;D^{-c/d}.$ (28)
Proof.
$\min_{j\neq i}l(g(Z_{i}),g(Z_{j}))$ $\leq\min_{j\neq i}\;L_{l}\;\cdot\;||Z_{i}-Z_{j}||$ (29)
$=L_{l}\cdot\min\{M,\min_{j\neq i}||Z_{i}-Z_{j}||\}$
[…]
By applying Theorem 1 we have:
$\mathbb{E}_{\mathcal{Z}_{D}}[\min_{j\neq i}l(g(Z_{i}),g(Z_{j}))]\leq L_{l}\;C\;D^{-c/d}$ (30)
for constant $C=C(d,L,\epsilon,M,C_{\epsilon})$, $c=c(L,\epsilon,M,C_{\epsilon})$ and large enough $D>D_{0}(d,L,\epsilon,M,C_{\epsilon})$, thus completing the proof. ∎
Meaning of a finite $\epsilon$-negative moment
  • Lebesgue–measure view
    Write $E_{t}:=\{z:q(z)\leq t\}$. Chebyshev gives
    $\text{Leb}(E_{t})\leq t^{-\epsilon}\!\int q^{1-\epsilon}=C_{\epsilon}\,t^{-\epsilon}.$ (31)
    Hence Assumption (A2) controls how large the very–low–density region can be; the smaller $\epsilon$, the larger that region may grow.
  • Rényi entropy view
    For order $\alpha>0$, the Rényi entropy is
    $H_{\alpha}(q)=-\frac{1}{\alpha-1}\log\int q^{\alpha}.$ (32)
    Setting $\alpha=1-\epsilon\in(0,1)$ (the Tsallis regime) and re-arranging,
    $\int q^{1-\epsilon}=e^{-(1-\epsilon)H_{1-\epsilon}(q)},$ (33)
    so finiteness of the $\epsilon$-negative moment is equivalent to finite sub-Rényi entropy of order $<1$. Smaller $\epsilon$ (order closer to $1$) corresponds to heavier low-density tails, which precisely slows the nearest-neighbour rate as captured in Theorem 1.
C.3 Derivation for Bayes Risk with Intrinsic Dimension Assumption
Theorem 3 (Bayes Risk and Context Length with Intrinsic Dimension Assumption).
Let $\mathcal{Z}$ be an intrinsic space satisfying Predictive Consistency and Uniform Information Gain, then the Bayes Risk $H(P,P_{l})$ of context length $l$ is Linear with respect to Intrinsic Dimension $dim(l)$. That is,
$H(P,P_{l})=-s\cdot dim(l)+Const$ (34)
Proof.
$H(P,P_{l})$ $=H(P)+D_{KL}(P,P_{l})$ (35)
$=H(P)+s\cdot(dim(\infty)-dim(l))$
$=-s\cdot dim(l)+Const$
An intuitive example for the ‘s-bits per dimension’ assumption: assuming that the vocab is an integer from 00 to $12^{dim(\infty)*s}-$ 1. assuming $P(x_{0}|x_{-\infty:0})=\delta_{x_{0},y}$, that is, the next token given $x_{-\infty:0}$ is sure to be $y$. $y$. For $P_{l}(x_{0}|x_{-\infty:0})$, the first $dim(l)*s$ digits of the integer (in binary representation) are known, but the remaining $(dim(\infty)-dim(l))*s$ digits are unknown, making a guess in these numbers yield $P_{l}(x_{0}|x_{-\infty:0})=1/2^{s*(dim(\infty)-dim(l))}$. Thus, $D_{KL,x_{0}}(P(x_{0}|x_{-\infty,0}),P_{l}(x_{0}|x_{-\infty,0}))=1*\log 1/(1/2^{s*(dim(\infty)-dim(l))})=s*(dim(\infty)-dim(l))$.
C.4 Derivation for Bayes Risk with Information Entropy Assumption
Theorem 4 (Bayes Risk and Context Length with Information Entropy Assumption).
Let $\mathcal{Z}$ be an intrinsic space satisfying Predictive Consistency and Linear Entropy Relationship of zero-tolerance limit, then the Bayes Risk $H(P,P_{l})$ of context length $l$ is Linear with respect to Intrinsic Entropy $H[q_{t}(\mathcal{Z})]$ where $q_{t}(\cdot)$ denotes probability density function. That is,
$H(P,P_{l})=-s\cdot H[q_{t}(Z)]+Const$ (36)
Proof.
[…]
$=H[p(x_{t}|x<t)]$
$=-s\cdot H[q_{t}(Z)]+b$
$=-s\cdot H[q_{t}(Z)]+Const$